OpenStudy (anonymous):

Using either the Echelon Method or the Gauss-Jordan Method, solve the following linear system. x + y + z = 2 2x + 3y -z = 5 x - y + 4z = 3

OpenStudy (anonymous):

1 1 1 2 2 3 -1 5 1 -1 4 3

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

matrix

OpenStudy (anonymous):

that's your coefficient matrix

OpenStudy (anonymous):

what do i multipy by for the matrix -3 ?

OpenStudy (anonymous):

it is hard to explain witout visualization.

OpenStudy (anonymous):

x=7/9 ... y=21/18 ... z=1/18

OpenStudy (anonymous):

i dont know how to write matrices on this ... but that is the result of matrix reduction ... i hope ... coulda made a mistake tho :S

OpenStudy (anonymous):

yes they work ... if only i could write up matrices i could show u the working

OpenStudy (anonymous):

do u want working out?